OpenStudy (anonymous):

It takes a large pipe 45 minutes to fill a pool. A smaller pipe can fill the pool in 110 minutes. Working together, how long does it take (to the nearest minute) for both pipes to fill the pool.

OpenStudy (amistre64):

how much of the pool is filled in 1 minute by each of them seperately?

OpenStudy (amistre64):

Lp = 1/45 parts per minute Sp = 1/110 parts per minute right?

OpenStudy (amistre64):

Lp + Sp working together fill the what part in one minute? 1/45 + 1/110 = 155/4950 is my easy math way :)

OpenStudy (amistre64):

155 parts out of 4950 are filled in 1 minute by both pipes working together

OpenStudy (amistre64):

31/990 might be an easier fraction to play with...

OpenStudy (anonymous):

so if the question was in hours would i just find how much of "x" is completed in one hour?

OpenStudy (amistre64):

Yep.... that would be it...then figure out how long it takes to do the entire "job"

OpenStudy (amistre64):

31 minutes and 29/31 left to fill ------ 31| 990 93 ---- 60 31 ---- 29

OpenStudy (amistre64):

id say 32 minutes at most :)

OpenStudy (anonymous):

you lost me

OpenStudy (anonymous):

i got how we had 155/4950

OpenStudy (anonymous):

then we simplified?

OpenStudy (amistre64):

if I did the addition right: we get 155 parts out of 4950 that are being doen in one minute 155 divides 4950 how many time?

OpenStudy (anonymous):

31?

OpenStudy (amistre64):

155 + 155 + 155 + 155....till you get near 4950

OpenStudy (anonymous):

OHHH

OpenStudy (anonymous):

ok i get it THANKS

OpenStudy (amistre64):

155 31 ----- 155 365 ----- 4805 right?

OpenStudy (anonymous):

oh 32

OpenStudy (amistre64):

4950 -4805 ------ 145 more parts to fill out of 155 per minute :)

OpenStudy (amistre64):

which is close to 1 minute so round to 32 :)

OpenStudy (anonymous):

instead of confusing yourself just think about it as if you were filling up your pool asnd you really wanted to go swim do you decide to use both of them so it can go faster one of you freinds asks you about what time will it be done and and you give her an amout of min. whatr would you tell him/her?

OpenStudy (amistre64):

that last part always used to get me... until I realized the the "155" took one full minute and we only needed "145" of those: 145 --- is the reminder of the minute which is close enough to 1 right :) 155

OpenStudy (anonymous):

ok that makes sense

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

who u talkin 2...

OpenStudy (anonymous):

not you

OpenStudy (anonymous):

i no but i want to know which person is right i am trying to figure it our myself and btw dnt be ruse witrh me u dnt know who u tlkkin to...